# Extending class group action attacks via sesquilinear pairings

### v1

@@ 8 When $N$ divides the discriminant
Let $E$ be an elliptic curve oriented by $\mathcal{O}=\mathbb{Z}[\tau]$. Let $\mathbb{F}$ be a finite field containing the $m$-th roots of unity. Suppose $E[m]=E[m](\mathbb{F})$ is a cyclic $\mathcal{O}$-module. Let $P\in E[m]$. Then the multiplicative order $T^{\prime}_{m}(P,P)$ is $m/t$ where $t$ is the minimal positive integer such that $[t]E[m]\subseteq\mathcal{O}P$.
@@
is non-degenerate and, for $P$ of order $m$, there exists a $Q$ so $t_{m}(P,Q)$ has order $m$ (Proposition 1). Let $t$ be the minimal positive integer for which $[t]E[m]\subseteq\mathcal{O}P$. Then $[t]Q=[a+\tau b]P$ for some $a,b$ having $\gcd(a,b,m)=1$.
@@
$=(t_{m}([\tau]P,P)^{a}t_{n}(P,P)^{N(\tau)b},t_{m}(P,P)^{a}t_{m}([\tau]P,P)^{b}).$
The left side has order $m/t$. Thus the right side has order $m/t$; applying a linear transformation, $T^{\prime}_{m}(P,P)$ must have order $m/t$.
@@
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented elliptic curves. Suppose there exists an oriented $\phi:E\rightarrow E^{\prime}$ of known degree $d$. Let $m$ be smooth, coprime to $d$, and chosen so that there are only polynomially many square roots of $1$ modulo $m$. Suppose there exists $\tau\in\mathcal{O}$ with norm $N(\tau)\equiv\Delta_{\mathcal{O}}\equiv 0\pmod{m}$. Suppose that $E[m]$ is a cyclic $\mathcal{O}$-module. Suppose $P\in E[m]$ such that $\mathcal{O}P=E[m]$. Then a subset $S\subset E[m]$ of polynomial size containing $\phi(P)$ can be computed in polynomially many operations in the field of definition of $E[m]$.
@@
As explained in [4, Section 5.2], we can find $\tau\in\mathcal{O}$ such that $N(\tau)\equiv Tr(\tau)\equiv 0\pmod{m}$. Hence the minimal polynomial of $\tau$ is $x^{2}$ modulo $m$, and $\tau^{2}$ acts as 00 on $\mathcal{O}/m\mathcal{O}$. Choose a point $P\in E[m]$ such that $\mathcal{O}P=E[m]$. Choose a point $P^{\prime}\in E^{\prime}[m]$ such that $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Then $T^{\prime}_{m}(P,P)$ and $T^{\prime}_{m}(P^{\prime},P^{\prime})$ have order $m$ by Proposition 2. Then
@@
Using a discrete logarithm, we can compute $N(\lambda)\pmod{m}$. Write $\lambda=a+b\tau$. Then $N(\lambda)\equiv a^{2}\pmod{m}$. Since the factorization of $m$ is known, by assumption, we have an efficiently computable set of polynomial size of possible values of $a$. Compute $[a][\tau]P^{\prime}$. For the correct $a$, this is the image of $[\tau]P$ under $\phi$, since
$\phi[\tau]P=\lambda[\tau]P^{\prime}=[a][\tau]P^{\prime}.$
Trying all possible values of $a$, we obtain the set $\{[a][\tau]P^{\prime}:a^{2}\equiv N(\lambda)\pmod{m}\}$ required by the statement.
The results of [4, 13] can now be applied if $m$ is a smooth square, $d$ is powersmooth and $m>d$, to reduce to a solvable instance of SIDH. Alternatively, loosening the restriction that $m$ is a square will be possible with the new generalizations of SIDH mentioned above [3].
@@
The following example is based on an example communicated to the authors by Wouter Castryck. Let $E:y^{2}=x^{3}+x$. Let $p$ be a prime of the form $14\cdot 3^{r}-$ 1 with $r>0$. This curve is supersingular with $j=1728$ and endomorphisms $[i]:(x,y)\mapsto(-x,iy)$ and $\pi_{p}:(x,y)\mapsto(x^{p},y^{p})$. Let
@@
Then $\tau^{2}=-\frac{p+1}{4}=-3^{r}$, so $N(\tau)=3^{r}$. Since $\pi_{p}^{2}=[-p]$, $E(\mathbb{F}_{p^{2}})\cong(\mathbb{Z}/(p+1)\mathbb{Z})^{2}\cong(\mathbb{Z}/4\cdot 3^{r}\mathbb{Z})^{2}$ [25, Ex. 5.16.d]. Therefore $E[3^{r}]\subseteq E(\mathbb{F}_{p^{2}})$.
We can choose $m=3^{r}$, which is a square when $r$ is even. Let $Q$ be an $\mathcal{O}$-generator of $E[3^{r}]$. Then by Proposition 2, $T^{\prime}_{m}(Q,Q)$ has order $3^{r}$, and the polynomially many operations to run the attack of Theorem 8.1 take place in $\mathbb{F}_{p^{2}}$. The SIDH portion of the attack requires that $d<m=3^{r}$.
By contrast, using the methods of [4, Section 6.1], one would need a generalized pairing value of $T^{\tau}_{m}$ for which only methods of computation taking place in the field of definition of $E[3^{2r}]$ are known.
@@ 9 Supersingular class group action in the presence of another orientation
Let $E$ and $E^{\prime}$ be elliptic curves for both of which we know orientations by two orders $\mathcal{O}$ and $\mathcal{O}^{\prime}$ which generate a rank $4$ sub-order of the endomorphism ring. Let $\phi:E\rightarrow E^{\prime}$ be an isogeny of known degree $d$. Let $m$ be smooth, coprime to the discriminants of $\mathcal{O}$ and $\mathcal{O}^{\prime}$, and suppose $1$ has only polynomially many square roots modulo $m$. Suppose $\phi$ respects both the $\mathcal{O}$ and $\mathcal{O}^{\prime}$ orientations. Let $P\in E[m]$. Then a subset of $E^{\prime}[m]$ of polynomial size containing $\phi P$ can be computed in polynomially many operations in the field of definition of $E[m]$.
@@
By assumption, $m$ is coprime to the discriminants. Thus we can choose $P\in E[m]$ so that $\mathcal{O}P=E[m]$. Similarly, let $P^{\prime}\in E^{\prime}[m]$ be chosen so that $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Suppose $\phi P=[\lambda]P^{\prime}$ for some $\lambda\in\mathcal{O}$. As in Theorem 6.1, we can compute $N(\lambda)$ modulo $m$.
Let $\sigma\in\mathcal{O}^{\prime}$ be chosen to have norm coprime to $m$ and be of the form $\sigma=a\sigma_{1}+b\sigma_{2}$ with $a$ and $b$ coprime to $m$, where $\sigma_{1}\in\mathcal{O}$ and $\sigma_{2}$ is perpendicular to $\mathcal{O}$ in the geometry afforded by the norm map on the quaternion algebra, so that $\sigma_{2}\lambda=\overline{\lambda}\sigma_{2}$. Write $\lambda^{\sigma}$ for an element which participates in the equivalence $\lambda^{\sigma}\sigma\equiv\sigma\lambda\pmod{m}$. This can be computed by taking $\lambda^{\sigma}=\sigma\lambda\overline{\sigma}N$ where $N$ is an integer equivalent to $N(\sigma)^{-1}\pmod{m}$. Then $\overline{\lambda^{\sigma}}\equiv a\overline{\lambda}+b{\lambda}\pmod{m}$. Then
@@
$=\widehat{T}_{m}^{\tau}(\sigma[\lambda]P^{\prime},[\lambda]P^{\prime})$
@@
Using knowledge of $\overline{\lambda^{\sigma}}\lambda\equiv aN(\lambda)+b\lambda^{2}$ and $N(\lambda)$, both modulo $m$, and recalling $b$ is coprime to $m$, we obtain $\lambda^{2}$ modulo $m$.
Using $N(\lambda)$ and $\lambda^{2}$, we can solve for polynomially many possibilities for $\lambda$ modulo $m$. Then we have obtained $\phi P$. From this we can compute any other $\phi R$ by solving for $R=[\mu]P$ and observing that $\phi R=\phi[\mu]P=[\mu]\phi P$.
@@ Acknowledgements
The authors are grateful to Damien Robert and Wouter Castryck for helpful discussions. Both authors are supported by NSF CAREER CNS-1652238 and NSF DMS-2401580 (PI K. E. Stange).

### v2

@@ 8 When $m$ divides the discriminant
Let $E$ be an elliptic curve oriented by $\mathcal{O}=\mathbb{Z}[\tau]$. Let $\mathbb{F}$ be a finite field containing the $m$-th roots of unity. Suppose $E[m]=E[m](\mathbb{F})$ is a cyclic $\mathcal{O}$-module. Let $P\in E[m]$ have order $m$. Then the multiplicative order $T^{\prime}_{m}(P,P)$ is at least $m/t$ where $t$ is the minimal positive integer such that $[t]E[m]\subseteq\mathcal{O}P$.
@@
is non-degenerate and, for $P$ of order $m$, there exists a $Q$ so $t_{m}(P,Q)$ has order $m$ (Proposition 1). Let $t$ be the minimal positive integer for which $[t]E[m]\subseteq\mathcal{O}P$. Then $[t]Q=[a+\tau b]P$. Using Proposition 1,
@@
$=\left(t_{m}([\tau]P,P)^{a}t_{m}([\tau]P,[\tau]P)^{b},t_{m}(P,P)^{a}t_{m}(P,[\tau]P)^{b}\right)$
$=\left(t_{m}([\tau]P,P)^{a}t_{m}(P,P)^{N(\tau)b},t_{m}(P,P)^{a+Tr(\tau)b}t_{m}([\tau]P,P)^{-b}\right).$
The left side has order $m/t$ by Proposition 1. Thus the right side has order $m/t$. This is the image of $T^{\prime}_{m}(P,P)$ via a linear transformation of determinant $N(\tau)b^{2}+a^{2}+Tr(\tau)ab=N(a+\tau b)$. Therefore $T^{\prime}_{m}(P,P)$ must have order at least $m/t$.∎
@@
Let $E$ and $E^{\prime}$ be $\mathcal{O}$-oriented elliptic curves. Suppose there exists an oriented isogeny $\phi:E\rightarrow E^{\prime}$ of known degree $d$. Let $m$ be smooth, coprime to $d$, and chosen so that there are only polynomially many square roots of $1$ modulo $m$. Suppose $m\mid\Delta_{\mathcal{O}}$. Suppose that $E[m]$ is a cyclic $\mathcal{O}$-module. Suppose $P\in E[m]$ such that $\mathcal{O}P=E[m]$, and $P^{\prime}\in E^{\prime}[m]$ such that $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Then there exists an efficiently computable point $Q\in E[m]$ of order $m$ such that a subset $S\subset E^{\prime}[m]$ of polynomial size containing $\phi(Q)$ can be computed in polynomially many operations in the field of definition of $E[m]$.
@@
Choose a point $P\in E[m]$ such that $\mathcal{O}P=E[m]$. Choose a point $P^{\prime}\in E^{\prime}[m]$ such that $\mathcal{O}P^{\prime}=E^{\prime}[m]$. Then $T^{\prime}_{m}(P,P)$ and $T^{\prime}_{m}(P^{\prime},P^{\prime})$ have order $m$ by Proposition 2. Then
@@
(Note that $\lambda$ is an endomorphism, so here we use compatibility with isogenies, not sesquilinearity in general.) Using a discrete logarithm, we can compute $N(\lambda)\pmod{m}$.
Observe that the definition of $T_{m}^{\prime}$ actually depends only on $\tau$ modulo $m$, and hence on $\mathbb{Z}[\tau]$ modulo $m$. So we now choose $\tau$ in a specific way, possibly not generating all of $\mathcal{O}$ but only generating it modulo $m$. In short, we claim the existence of $\tau\in\mathcal{O}$ with certain properties, namely that
  1. 1.
    $\mathbb{Z}[\tau]\equiv\mathcal{O}$ modulo $m$;
  2. 2.
    $Tr(\tau)\equiv N(\tau)\equiv 0\pmod{m^{\prime}}$ where $m^{\prime}=m/4$ if $4\mid m$; $m^{\prime}=m/2$ if $m\equiv 2\pmod{4}$; and $m^{\prime}=m$ otherwise.
The existence of such a $\tau$ is a consequence of $m\mid\Delta_{\mathcal{O}}$, as follows. A generator for $\mathcal{O}$ is given by $\sigma=\frac{\Delta+\sqrt{\Delta}}{2}$ having trace $\Delta$ and norm $\frac{1}{4}(\Delta-\Delta^{2})$. Then $\tau=2\sigma$ already has the required properties if $m$ is odd. If $m$ is even, then $4\mid\Delta$, and the norm is divisible by $m^{\prime}$, so $\tau=\sigma$ suffices. Hence the minimal polynomial of $\tau$ is $x^{2}$ modulo $m^{\prime}$, and $[\tau^{2}]E^{\prime}[m]\subseteq E^{\prime}[m/m^{\prime}]\subseteq E^{\prime}[4]$.
Write $\lambda\equiv a+b\tau$ modulo $m$. Then $N(\lambda)\equiv a^{2}\pmod{m^{\prime}}$. Since the factorization of $m^{\prime}$ is known, by assumption, we have an efficiently computable set of polynomial size of possible values of $a$. Compute $[a][\tau]P^{\prime}$. For the correct $a$, this is the image of $[\tau]P$ under $\phi$ up to addition of a $4$-torsion point, since
$\phi[\tau]P=[\lambda][\tau]P^{\prime}\in[a][\tau]P^{\prime}+E^{\prime}[4].$
Trying all possible values of $a$, and setting $Q=[\tau]P$, we obtain the set
$\{[a][\tau]P^{\prime}:a^{2}\equiv N(\lambda)\pmod{m}\}+E^{\prime}[4]$
required by the statement. Observe that $P,Q$ form a basis for $E[m]$ by construction, so $Q$ has order $m$. ∎
For each possible value of $a$, we have a putative $\phi([\tau]P)$, i.e. the action of $\phi$ on a single $m$-torsion point. The results of [CHM+23, FFP24] can now be applied if $m$ is a smooth square, $d$ is powersmooth and $m>4d$ , to reduce to SIDH. Alternatively, loosening the restriction that $m$ is a square will be possible with the new generalizations of SIDH mentioned above [CDM+24].
The following example demonstrates a new growing family of parameters for which solving the class group action problem (with known degree) is polynomial instead of exponential using Theorem 8.1.
@@
This is based on an example communicated to the authors by Wouter Castryck. Let $E:y^{2}=x^{3}+x$. Let $p$ be a prime of the form $14\cdot 3^{r}-$ 1 with $r>0$. This curve is supersingular with $j=1728$ and endomorphisms $[i]:(x,y)\mapsto(-x,iy)$ and $\pi_{p}:(x,y)\mapsto(x^{p},y^{p})$. Let
@@
Then $\tau^{2}=-\frac{p+1}{4}=-3^{r}$, so $N(\tau)=3^{r}$ and $Tr(\tau)=0$. Let $\mathcal{O}=\mathbb{Z}[\tau]$, having $N(\tau)\mid\Delta_{\mathcal{O}}$. Let $m=3^{r}$. Then $m\mid\Delta_{\mathcal{O}}$.
Since $\pi_{p}^{2}=[-p]$, $E(\mathbb{F}_{p^{2}})\cong(\mathbb{Z}/(p+1)\mathbb{Z})^{2}\cong(\mathbb{Z}/4\cdot 3^{r}\mathbb{Z})^{2}$ [Sil09, Ex. 5.16.d]. Therefore $E[3^{r}]\subseteq E(\mathbb{F}_{p^{2}})$.
Let $Q$ be an $\mathcal{O}$-generator of $E[3^{r}]$. Then by Proposition 2, $T^{\prime}_{m}(Q,Q)$ has order $3^{r}$, and the polynomially many operations to run the attack of Theorem 8.1 take place in $\mathbb{F}_{p^{2}}$. The SIDH portion of the attack requires that $4d<m=3^{r}$.
By contrast, using the methods of [CHM+23, Section 6.1], one would need a generalized pairing value for which only methods of computation taking place in the field of definition of $E[3^{2r}]$ are known, and this field degree grows exponentially with $r$. (Specifically, in [CHM+23, Section 5, p. 20], the authors give an estimated runtime for the pairing needed in the attack, noting that their method requires dividing a point by $m$ and working in the resulting field extension of degree as much as $O(m^{2})$.) That means that what was an exponential runtime in terms of $r$ under [CHM+23] becomes a polynomial one using Theorem 8.1.
@@ 9 Supersingular class group action in the presence of another orientation
The following theorem shows that, if we have two distinct orientations respected by $\phi$, then we can recover the action of $\phi$.
Suppose $\mathcal{O}\subseteq\operatorname{End}(E)$. We use the notation $\mathcal{O}^{\perp}$ for the quadratic order orthogonal to $\mathcal{O}$ within the endomorphism ring, with respect to the geometry induced by the quaternion norm.
@@
Let $E$ and $E^{\prime}$ be supersingular elliptic curves for both of which we know orientations by two quadratic orders $\mathcal{O}$ and $\mathcal{O}^{\prime}$ which together generate a rank $4$ sub-order of the endomorphism ring. Let $\phi:E\rightarrow E^{\prime}$ be an isogeny of known degree $d$. Let $m$ be smooth, coprime to the discriminants of $\mathcal{O}$ and $\mathcal{O}^{\prime}$, and suppose $1$ has only polynomially many square roots modulo $m$. Suppose $\phi$ respects both the $\mathcal{O}$ and $\mathcal{O}^{\prime}$ orientations. Suppose $\mathcal{O}$-module generators are known for both $E[m]$ and $E^{\prime}[m]$. Suppose, finally, that $\mathcal{O}^{\perp}$ has elements of norm coprime to $m$. Let $P\in E[m]$. Then a subset of $E^{\prime}[m]$ of polynomial size containing $\phi P$ can be computed in a polynomial number of operations in the field of definition of $E[m]$.
@@
Suppose $E[m]=\mathcal{O}P$ and $E^{\prime}[m]=\mathcal{O}P^{\prime}$.
Let $\sigma\in\operatorname{End}(E)$ be chosen to have norm coprime to $m$. Write $\lambda^{\sigma}$ for an element which participates in the equivalence $\lambda^{\sigma}\sigma\equiv\sigma\lambda\pmod{m}$. Suppose $\lambda^{\sigma}\in\mathcal{O}$. Then
@@
$=\widehat{T}_{m}^{\tau}([\sigma][\lambda]P^{\prime},[\lambda]P^{\prime})$
@@
We will now apply the above for two specially chosen $\sigma\in\operatorname{End}(E)$. Since we can compute the action of $\mathcal{O}$ and $\mathcal{O}^{\prime}$, we can compute the action of anything they generate. Thus, we choose $\sigma_{1}=1$ (so $\lambda^{\sigma_{1}}=\lambda$), and then some $\sigma_{2}\in\mathcal{O}^{\perp}$, so $\lambda^{\sigma_{2}}=\overline{\lambda}$. Assuming that $\mathcal{O}^{\perp}$ contains elements of norm coprime to $m$, from this, we obtain both $N(\lambda)$ and $\lambda^{2}$ modulo $m$.
Using $N(\lambda)$ and $\lambda^{2}$, we can solve for polynomially many possibilities for $\lambda$ modulo $m$ (this requires smoothness, so $m$ can be factored). Then we have obtained $\phi P$. From this we can compute any other $\phi R$ by solving for $R=[\mu]P$ and observing that $\phi R=\phi[\mu]P=[\mu]\phi P$.∎
@@ Acknowledgements
The authors are grateful to Wouter Castryck, Steven Galbraith, Marc Houben, Massimo Ostuzzi, Lorenz Panny, James Rickards and Damien Robert for helpful discussions. They are also grateful to the anonymous referees for feedback which much improved the paper. Both authors have been supported by NSF CAREER CNS-1652238 and NSF DMS-2401580 (PI K. E. Stange).